OpenStudy (anonymous):

write an equation of the line containing the points (3,2) and (0,-3) us the form y=mx+b plzzzz help meee asap

OpenStudy (anonymous):

a^

OpenStudy (anonymous):

b^

OpenStudy (anonymous):

c^

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

(x1,y1) = (3,2) (x2,y2) = (0,-3) use the slope formula \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\] to find the slope

OpenStudy (anonymous):

i was never taught this though

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

subtract the y coordinates subtract the x coordinates make sure you subtract in the same order then divide the two differences (y diff over x diff)

OpenStudy (anonymous):

what? i dont get it @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what are the y coordinates of these two points (3,2) and (0,-3)

OpenStudy (anonymous):

ive never done that before

jimthompson5910 (jim_thompson5910):

hmm perhaps you're in a lesson that the teacher hasn't gotten to yet?

OpenStudy (anonymous):

its a scantron upf\date and ive never been taught that. my teaccher said to find help on the internet. are you supposed to learn this is 6th gradr?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

im just going to guess @jim_thompson5910

jimthompson5910 (jim_thompson5910):

if the teacher didn't get to this part, then I don't see how it's fair she expects you to know this material

OpenStudy (anonymous):

b ?

jimthompson5910 (jim_thompson5910):

anyways, the y coordinates of (3,2) and (0,-3) are 2 and -3, agreed?

OpenStudy (anonymous):

yes i think

OpenStudy (anonymous):

is it b

jimthompson5910 (jim_thompson5910):

subtract the y coordinates 2 minus (-3) = 2 - (-3) = 2+3 = 5

jimthompson5910 (jim_thompson5910):

what are the x coordinates of the two points?

OpenStudy (anonymous):

dont know im going to guess